# Flow-GRPO: Training Flow Matching Models via Online RL

### v3

@@ 5.1 Experimental Setup
Table 2: Performance on Compositional Image Generation, Visual Text Rendering, and Human Preference benchmarks, evaluated by task performance, image quality, and preference scores. Acc: Accuracy; ImgRwd: ImageReward; UniRwd: UnifiedReward.
@@
Flow-GRPO (w/o KL) 23.41 6.15 4.16 1.24 3.57
Flow-GRPO (w/ KL) 23.31 5.92 4.22 1.28 3.66
@@ 5.2 Main Results
Figure 1 and Table 1 show Flow-GRPO’s GenEval performance steadily improving during training, ultimately outperforming GPT-4o. This occurs while maintaining both image quality metrics and preference scores on DrawBench, a benchmark with diverse and comprehensive prompts for evaluating general model capabilities. Figure 3 offers qualitative comparisons. Beyond Compositional Image Generation, Table 2 details evaluations on Visual Text Rendering and Human Preference tasks. Flow-GRPO improved text rendering ability, again without decreasing image quality metrics and preference scores on DrawBench. See Figures 9, 10 & 11 in Appendix C.4 for related qualitative examples. For the Human Preference task, image quality did not decrease without KL regularization. However, we found that omitting KL caused a collapse in visual diversity, a form of reward hacking discussed further in Section 5.3. These results demonstrate that Flow-GRPO boosts desired capabilities while causing very little degradation to image quality or visual diversity.
@@ B.3 Hyperparameters Specification
Except for $\beta$, GRPO hyperparameters are fixed across tasks. We use a sampling timestep $T=10$ and an evaluation timestep $T=40$. Other settings include a group size $G=24$, an noise level $a=0.7$ and an image resolution of 512. The KL ratio $\beta$ is set to 0.004 for GenEval and Text Rendering, and 0.001 for Pickscore. We use Lora with $\alpha=64$ and $r=32$.
@@ Appendix C Extended Experimental Results
Offline variants use a fixed pretrained model for data collection, while online variants update their data collection model every 40 steps. As shown in Figure 6, Flow-GRPO outperforms all other methods. The figure also indicates that DPO and SFT improve over time. In contrast, RWR does not, which aligns with experimental findings on RWR in [12]. Additionally, Online DPO surpasses offline DPO, aligning with [15]’s finding that online DPO performs better. For the second-best online DPO, a hyperparameter search on its key parameter $\beta$ revealed that smaller values are not always optimal; excessively small $\beta$ values can cause training collapse.
Figure 6: Comparison of Flow-GRPO and Other Alignment Methods.
@@ C.2 Effect of Denoising Reduction
(a). Visual Text Rendering
(b). Human Preference Alignment
@@ C.3 Learning Curves with or without KL
(a). Compositional Image Generation
(b). Visual Text Rendering
(c). Human Preference Alignment
@@ C.4 Additional Qualitative Results
Figures 9, 10 & 11 qualitatively compare SD3.5-M with its Flow-GRPO enhanced versions (with and without KL regularization) using GenEval, OCR and PickScore rewards, respectively. Flow-GRPO with KL regularization improves the target capability while maintaining image quality and minimizing reward-hacking. Conversely, removing the KL constraint significantly degrades image quality and diversity.

### v4

@@ 5.1 Experimental Setup
Table 2: Performance on Compositional Image Generation, Visual Text Rendering, and Human Preference benchmarks, evaluated by task performance on test prompts, and by image quality and preference scores on DrawBench prompts. ImgRwd: ImageReward; UniRwd: UnifiedReward.
@@
Flow-GRPO (w/o KL) 23.41 6.15 4.16 1.24 23.56 3.57
Flow-GRPO (w/ KL) 23.31 5.92 4.22 1.28 23.53 3.66
@@ 5.2 Main Results
Figure 1 and Table 1 show Flow-GRPO’s GenEval performance steadily improving during training, ultimately outperforming GPT-4o. This occurs while maintaining both image quality metrics and preference scores on DrawBench, a benchmark with diverse and comprehensive prompts for evaluating general model capabilities. Figure 3 offers qualitative comparisons. Beyond Compositional Image Generation, Table 2 details evaluations on Visual Text Rendering and Human Preference tasks. Flow-GRPO improved text rendering ability, again without decreasing image quality metrics and preference scores on DrawBench. See Figures 10, 11 & 12 in Appendix C.4 for related qualitative examples. For the Human Preference task, image quality did not decrease without KL regularization. However, we found that omitting KL caused a collapse in visual diversity, a form of reward hacking discussed further in Section 5.3. These results demonstrate that Flow-GRPO boosts desired capabilities while causing very little degradation to image quality or visual diversity.
@@ 6 Conclusion
Acknowledgements
We would like to thank Mingwu Zheng for discussions on the proof and Zhanhui Zhou for suggestions on the writing.
@@ B.3 Hyperparameters Specification
Except for $\beta$, GRPO hyperparameters are fixed across tasks. We use a sampling timestep $T=10$ and an evaluation timestep $T=40$. Other settings include a group size $G=24$, an noise level $a=0.7$ and an image resolution of 512. The KL ratio $\beta$ is set to 0.04 for GenEval and Text Rendering, and 0.01 for Pickscore. We use Lora with $\alpha=64$ and $r=32$.
@@ Appendix C Extended Experimental Results
Offline variants use a fixed pretrained model for data collection, while online variants update their data collection model every 40 steps. As shown in Figure 6, Flow-GRPO outperforms all other methods. The figure also indicates that DPO and SFT improve over time. In contrast, RWR does not, which aligns with experimental findings on RWR in [12]. Additionally, Online DPO surpasses offline DPO, aligning with [15]’s finding that online DPO performs better. For the second-best online DPO, a hyperparameter search on its key parameter $\beta$ revealed that smaller values are not always optimal; excessively small $\beta$ values can cause training collapse. Figure 7 shows similar conclusions.
Figure 6: Comparison of Flow-GRPO and Other Alignment Methods on the Human Preference Alignment task.
Figure 7: Comparison of Flow-GRPO and Other Alignment Methods on the Compositional Image Generation task.
@@ C.2 Effect of Denoising Reduction
(a) Visual Text Rendering
(b) Human Preference Alignment
@@ C.3 Learning Curves with or without KL
(a) Compositional Image Generation
(b) Visual Text Rendering
(c) Human Preference Alignment
@@ C.4 Additional Qualitative Results
Figures 10, 11 & 12 qualitatively compare SD3.5-M with its Flow-GRPO enhanced versions (with and without KL regularization) using GenEval, OCR and PickScore rewards, respectively. Flow-GRPO with KL regularization improves the target capability while maintaining image quality and minimizing reward-hacking. Conversely, removing the KL constraint significantly degrades image quality and diversity.